OpenStudy (anonymous):

Let f(x)=e^−x^2. (a) What horizontal or vertical asymptotes does f(x) have? (b) Find the intervals of increase or decrease. (c) Find the local maximum and minimum values. (d) Where is f(x) concave downward? (e) Where are the inflection points?

OpenStudy (anonymous):

can you write the eqaution out with the drawing tool.. im sorry but i think it would make the equation much more clear

OpenStudy (anonymous):

\[f(x)=e^{-x^2}\] like this ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

kk ! now tell me can you find \[\lim_{x \rightarrow \pm \infty} f(x)\] ?!

OpenStudy (anonymous):

I believe it is 0 for both

OpenStudy (anonymous):

good ! So is it finite ?! ( just answer) what can you say about the asymptotes of "f" ?!

OpenStudy (anonymous):

the horizontal asymptote is 0 and so is the vertical

OpenStudy (anonymous):

there is no vertical asymptote ! ! \[\lim_{x \rightarrow \pm \infty}f(x)=a \] finite ! y=a is an horizontal asymptote ! so y=0 is the horizontal asymptote ! now ! if \[\lim_{x \rightarrow a}f(x)= \pm \] ! then you can say x=a is a verttical asymptote ! now ! we agree that y=0 is the only asymptote of 'f' is that clear ?!

OpenStudy (anonymous):

**\[\lim_{x \rightarrow a}f(x)= \pm \infty \]

OpenStudy (anonymous):

ok that makes sense

OpenStudy (anonymous):

@fheard still there ?!

OpenStudy (anonymous):

yeah but then how do I find increase or decrease

OpenStudy (anonymous):

you can't put -infinity,a and a,infinity can you?

OpenStudy (anonymous):

wait a is 0 right

OpenStudy (anonymous):

yes ! you're right ! okk can you calculate f' (the derviative) ...and see where It's negative and positive kk ?!

OpenStudy (anonymous):

still there ! ?

OpenStudy (anonymous):

yeah the derivative is -2x*e^-x^2

OpenStudy (anonymous):

alright I got it now thanks for the help.

OpenStudy (anonymous):

welcome To Openstudy !